Evaluate
f(-2/3) if f(x) = x/2

Answer by checkley77(12569)

f(-2/3)
f(x) = x/2
f(-2/3)=(-2/3)/2
f(-2/3)=-2/3*1/2
f(-2/3)=-2/6
f(-2/3)=-1/3 answer.